OpenStudy (anonymous):

A game has 40 chips, each labeled with a number from 1 through 40. Ben randomly chooses one chip. What is the probability that he chooses 10 or a multiple of 8? A.1/8 B.3/20 C.1/6 D.1/320

OpenStudy (anonymous):

B? @peachpi

OpenStudy (anonymous):

@peachpi @LynFran

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5/40 + 1/40 = 6/40 = 3/20

OpenStudy (anonymous):

yeahh. :) thanks @peachpi